OpenStudy (anonymous):

find the vertex of y=x^2-14x+49

OpenStudy (cwrw238):

convert x^2 - 14x into the form (x - a)^2 - a^2 where a = 14/ 2 then add the 49

OpenStudy (cwrw238):

the last number will give you the y coordinate of the vertex

OpenStudy (anonymous):

I did. im still getting the incorrect answer

OpenStudy (cwrw238):

(x - 7)^2 - 49 + 49

OpenStudy (cwrw238):

= (x -7)^2 + 0 the vertex is at (7,0)

OpenStudy (anonymous):

oh wow ! I figure that was it! I did something wrong but that is one of my multiple choice answer. thanks !!

OpenStudy (cwrw238):

thats ok - the x-coordinate is the value of a in (x -a)^2

OpenStudy (anonymous):

ok I see now. thanks